The following information has been gathered on the costs and effectiveness of the two treatments, A and B. In this problem, costs and consequences are not discounted.

Treatment A

Treatment B

Mortality rate

2%

5%

Life expectancy for survivors

30 years

20 years

Initial treatment cost

$20,000

$10,000

Follow up costs, year 1

$5,000

$1,000

Annual follow up costs, all subsequent years

$1,000

$500

  1. Calculate the incremental cost and incremental benefit of the treatment alternatives.

DC =

DE =

  1. What is the ICER?

ICER = DC/DE =
Answer #1

DC means the difference in the cost of the two treatments.

So we calculate the cost of treatment A first

Cost = 20000 + 5000 + 1000* 29 (because life expectancy is 30 years)

= 54000

Cost of treatment B = 10000+ 1000 + 500*19

= 20500

DC= 54000 - 20500 = 33500 $

DE stands for difference in effectiveness .

DE = 5-2 = 3

ICER = DC/DE = 33500/3 = 11166.7 $ per percent

So ICER is 11166.7 $ per percent mortality .

So it means treatment A decreases mortality by 1% for 11166.7$ extra cost.